OpenStudy (anonymous):

Geoff estimates that 27% of 121 is about 30. Which of the following statements is true? His estimate is way too high. His estimate is way too low. His estimate is good. There is not enough information.

OpenStudy (anonymous):

27% of 121 is 0.27 * 121. That's the exact value of it. See how close 30 is to that.

OpenStudy (anonymous):

27/100 * 121 = 30 27 * 121/100 = 27 * 1.21 = 32.67, so guess is close

OpenStudy (anonymous):

27/100 * 121 is not equal to 30.

OpenStudy (anonymous):

yes, but the estimate is good as it is so near 30

OpenStudy (anonymous):

ohhh, sorry,

OpenStudy (anonymous):

Yes, it's a good estimate. But 27 / 100 * 121 is not equal to 30.

OpenStudy (anonymous):

yeah, got that, was written in hurry, lolzz

OpenStudy (anonymous):

actually to find the percentage of 121 that 30 is, you should do: \[\frac{x}{100} * 121 = 30 iff x = \frac{30}{121}* 100 = 24.7933... \]

OpenStudy (anonymous):

so 27% is an overestimation by about 2.2%

OpenStudy (anonymous):

hmm, yeah u r right :)